# Yesterday’s News: Benchmarking Multi-Dimensional Out-of-Distribution Generalisation of Misinformation Detection Models

### v1

@@ Abstract
This paper introduces misinfo-general, a benchmark dataset for evaluating misinformation models’ ability to perform out-of-distribution generalisation. Misinformation changes rapidly, much quicker than moderators can annotate at scale, resulting in a shift between the training and inference data distributions. As a result, misinformation models need to be able to perform out-of-distribution generalisation, an understudied problem in existing datasets. We identify 6 axes of generalisation—time, event, topic, publisher, political bias, misinformation type—and design evaluation procedures for each. We also analyse some baseline models, highlighting how these fail important desiderata. We make this dataset and accompanying code publicly available [^1] https://github.com/ioverho/misinfo-general
[…]
booktabs
Yesterday’s News: Benchmarking Multi-Dimensional
Out-of-Distribution Generalisation of Misinformation Detection Models
Ivo Verhoeven, Pushkar Mishra, Ekaterina Shutova $\dagger$ ILLC, University of Amsterdam, The Netherlands $\ddagger$ Meta AI, London, United Kingdom i.o.verhoeven@uva.nl, pm576@cantab.ac.uk, e.shutova@uva.nl
@@ 1 Introduction
The field of misinformation detection aims to develop classification models that can reliably moderate online content. Despite burgeoning academic interest Wu et al. (2019); Zhou and Zafarani (2020) and impressive classification results on existing datasets, content with the explicit intent to deceive continues to propagate and cause harm.
The rapid evolution of online content, which significantly outpaces model development cycles, partially explains this. For example, recent emerging news events Nair (2024); Reuters (2024), containing novel entities or entity relations, spawned a deluge of misinformation resulting in detrimental societal effects. Given that domain expert labels are prohibitively expensive to procure, in such cases training datasets will lag behind the news landscape. As a result, the inference-time text distribution will quickly shift away from the training data. For misinformation detection to be successful during deployment, models need to be robust to distribution shifts—to mitigate future harms, models must leverage general rules learned from yesterday’s news.
Currently, this property of Out-of-Distribution (OoD) generalisation is lacking in many SoTA models. Performance degrades when evaluated on unseen time periods Bozarth and Budak (2020); Horne et al. (2020b); Kochkina et al. (2023); Stepanova and Ross (2023), events Lee et al. (2021); Ding et al. (2022), domains Hoy and Koulouri (2022); Kochkina et al. (2023); Verhoeven et al. (2024), cultures, or languages Horne et al. (2020a); Chu et al. (2021); Ozcelik et al. (2023).
@@
We partially attribute this to the state of existing datasets. While plentiful, these are usually small, have been collected over a short time-span, are centred around specific events, or contain a homogeneous set of publishers. To learn generalisation, however, modern NLP models require large, diverse pre-training datasets, especially when text or labels are noisy.
This paper presents a novel benchmark dataset that enables training and evaluation for OoD generalisation in misinformation detectors. It balances article quantity and publisher diversity with labelling quality by leveraging rich publisher-level metadata annotations. We identify six axes along which we test generalisation performance, representing the most comprehensive evaluation at scale: 1) time, 2) specific events, 3) topics, 4) publisher style, 5) political bias and 6) misinformation type.
@@
We train a baseline model and provide in-depth analysis of its performance to showcase this dataset’s features. We show generalisation to different classes of publishers to be particularly challenging. Further analysis using the available publisher-level metadata highlights some undesirable model behaviours, such as worse performance on less frequent publishers, and worse classification on different political biases. Finally, we find some initial evidence that the scale and diversity of this dataset can benefit generalisation ability.
@@ 2 Related Work
2.1 Generalisable Misinformation Detection
@@
Generalisation abilities of misinformation classifiers have been tested at smaller scales. Horne et al. (2020b) found that performance degrades quickly when evaluating on future events, which Bozarth and Budak (2020) corroborate and extend to changes in domain. The same issues have also been reported in misinformation detection in other modalities Stepanova and Ross (2023); Verhoeven et al. (2024).
@@
Results on existing benchmark datasets are generally not indicative of down-stream performance. Kochkina et al. (2023) found that performance within one dataset vastly overestimates performance on other datasets or time spans. Even when controlling for the time period or topic, Hoy and Koulouri (2022) found that models overfit to the training dataset and perform worse on similar but unseen datasets. In a recent systematic review of the literature, Xiao and Mayer (2024) come to the conclusion that “…detection tasks are often meaningfully distinct from the challenges that online services actually face. Datasets and model evaluation are often non-representative of real-world contexts, and evaluation frequently is not independent of model training” (p. 1).
@@
There has been some interest in developing generalisable or adaptive misinformation detection techniques. This has been attempted through multitask training Lee et al. (2021), utilising external agents Ding et al. (2022); Mosallanezhad et al. (2022), data resampling or active learning Hu et al. (2023), adversarial learning Lin et al. (2022), or gradient-based meta-learning Yue et al. (2022). While these research directions are promising, their utility for out-of-distribution misinformation detection has not been sufficiently tested on large, diverse benchmark data.
2.2 Misinformation Datasets
There exist more misinformation datasets than we can exhaustively list here. Instead, we highlight the $4$ predominant annotation schemes, from most fine-grained to most coarse-grained.
@@ 2.2 Misinformation Datasets
    Span: experts manually fact-check individual (but complete) statements Wang (2017); Martino et al. (2020); Poldvere et al. (2023)
@@
    Document: experts manually check the overall veracity of entire documents Potthast et al. (2017); Shu et al. (2019)
@@
    Claim: experts verify claims related to specific events, while article labels are generated by proxy from their stance to these claims Nielsen and McConville (2022)

### v2

# Yesterday’s News: Benchmarking Multi-Dimensional Out-of-Distribution Generalization of Misinformation Detection Models

@@ Abstract
This article introduces misinfo-general, a benchmark dataset for evaluating misinformation models’ ability to perform out-of-distribution generalization. Misinformation changes rapidly, much more quickly than moderators can annotate at scale, resulting in a shift between the training and inference data distributions. As a result, misinformation detectors need to be able to perform out-of-distribution generalization, an attribute they currently lack. Our benchmark uses distant labelling to enable simulating covariate shifts in misinformation content. We identify time, event, topic, publisher, political bias, misinformation type as important axes for generalization, and we evaluate a common class of baseline models on each. Using article metadata, we show how this model fails desiderata, which is not necessarily obvious from classification metrics. Finally, we analyze properties of the data to ensure limited presence of modelling shortcuts. We make the dataset and accompanying code publicly available [^1] https://github.com/ioverho/misinfo-general.
@@ 1 Introduction
The field of misinformation detection aims to develop classification models that can reliably moderate online content. Despite burgeoning academic interest Wu et al. (2019); Zhou and Zafarani (2020) and impressive classification results on existing datasets, mis- and disinformation content continues to propagate online and cause significant societal harm. The rapid evolution of online content, which significantly outpaces model development cycles, partially explains this. News, and more generally all online content, is valued primarily for its novelty. It will often contain unseen entities, events and entity-event relationships. Verifying this content manually using domain experts is prohibitively expensive, and typically requires context not available when news is emerging.
The misinformation datasets used for misinformation detector training, however, are collections of yesterday’s news, containing content collected well after the fact. Furthermore, these datasets typically have a narrow focus on particular events or misinformation forms. As a result, state-of-the-art moderation systems lag behind the news landscape, and encounter inference-time data distributions that have shifted away from its training data distribution. For misinformation detection to be successful at mitigating harms during deployment, and especially in situations with limited social or historical context (i.e., when news is emerging), models will need to be robust to distribution shifts.
Currently, this property of Out-of-Distribution (OoD) generalization is lacking in many SoTA NLP models, and this is especially true for misinformation detection models. Performance significantly degrades when evaluated on unseen:
  • time periods Bozarth and Budak (2020); Horne, Nørregaard, and Adali (2020); Kochkina et al. (2023); Stepanova and Ross (2023),
  • publishers Rashkin et al. (2017); Zhou et al. (2021),
  • events Lee et al. (2021); Cheng, Nazarian, and Bogdan (2021); Ding et al. (2022); Wu and Hooi (2022),
  • domains Hoy and Koulouri (2022); Kochkina et al. (2023); Verhoeven et al. (2024),
  • cultures, or languages Horne, Gruppi, and Adalı (2020); Chu, Xie, and Wang (2021); Ozcelik et al. (2023).
@@
We primarily attribute this to the present state of misinformation datasets. While plentiful, these are often small, collected over a short time span, centered around specific events, biased towards popular content, or contain a homogeneous set of publishers. These properties are generally believed to be detrimental to the generalization capabilities of modern NLP models, which require large, diverse pre-training datasets, especially when text or labels are noisy.
Creating a dataset that does enforce OoD generalization, however, is not easy. Given the expense involved in collecting these datasets, prior attempts at doing so have invariably had to make trade-offs in size, diversity, or label fidelity (see Section 3). As a result, these datasets are not representative of the misinformation landscape, and evaluation with such datasets will overestimate model performance during deployment Aïmeur, Amri, and Brassard (2023); Xiao and Mayer (2024); Kuntur et al. (2024)). Generating high-quality misinformation labels for a realistically sized, naturalistic dataset remains intractable due to the cost of domain experts and the inherent subjectivity present in online content. This article will not solve this problem. Instead, we focus on more accurately estimating the generalization gap.
Specifically, we present misinfo-general, a dataset meant for testing the generalization performance of automated misinformation detectors holistically. We do so by processing a distantly labelled series of corpora intended for publisher reliability labelling. While this introduces noise into the labels, we argue that the scale and diversity of the data make it useful for generalizability evaluation. To mitigate said noise, we perform extensive pre-processing of the data (Section 4), and post-hoc testing of dataset properties (Section 8). This should ensure a balance of article quantity and label quality, providing one with rich publisher-level metadata, across a long time-span, covering a multitude of events and topics.
@@
To showcase the utility of such a benchmark, we identify and operationalize six generalization axes—(1) time, (2) specific events, (3) topics, (4) publisher style, (5) political bias and (6) misinformation type. We then train a simple, yet representative baseline model. We find that generalization to different classes of publishers is particularly challenging, whereas within-publisher variation across years is smaller than expected. Using the metadata available to us, we provide additional analysis of publisher-level determinants of performance, and find some undesirable model behaviors not discussed in prior literature: less frequent publishers see degraded performance, and models treat different political biases differently. Juxtaposed to these results, we also find some initial evidence that the scale and diversity of this dataset can benefit model generalization ability when trained on.
@@ 2 Related Work
2.1 Generalizable Misinformation Detection
@@
Generalization abilities of misinformation classifiers have been tested in many settings, at smaller scales. Horne, Nørregaard, and Adali (2020) found that performance degrades quickly when evaluating on future events, which Bozarth and Budak (2020) corroborate and extend to changes in domain. The same issues have also been reported in misinformation detection in other modalities Stepanova and Ross (2023); Verhoeven et al. (2024). Zhou et al. (2021) find that models tend to overfit to publisher idiosyncrasies more than article content, especially in publisher-level annotated datasets.
@@
Results on existing benchmark datasets are generally not indicative of downstream performance. Kochkina et al. (2023) found that performance within one dataset vastly overestimates performance on other datasets or time spans. Even when controlling for the time period or topic, Hoy and Koulouri (2022) found that models overfit to the training dataset and perform worse on similar but unseen datasets. In a recent systematic review of the literature, Xiao and Mayer (2024) come to the conclusion that:
…detection tasks are often meaningfully distinct from the challenges that online services actually face. Datasets and model evaluation are often non-representative of real-world contexts, and evaluation frequently is not independent of model training (p. 1)
This sentiment matches the earlier discussion in Aïmeur, Amri, and Brassard (2023); current misinformation benchmarks and evaluation setups can yield deceptively high performance scores.
@@
Despite this paucity in benchmarks and labels, there has been some interest in developing generalizable or adaptive misinformation detection techniques. This has been attempted through weak supervision Shu et al. (2020), multitask training Lee et al. (2021), utilizing external agents Ding et al. (2022); Mosallanezhad et al. (2022), data resampling or active learning Hu et al. (2023), adversarial learning Lin et al. (2022), or gradient-based meta-learning Zhang et al. (2021a); Yue et al. (2022). While these research directions are promising, their utility for out-of-distribution misinformation detection has not been sufficiently tested on large, diverse benchmark data.
2.2 Publisher Reliability Estimation
A related field to misinformation classification, especially when utilizing publisher-level labels, is publisher reliability estimation. Instead of yielding article level moderation decisions, a publisher reliability model uses the content of one or many articles from one publisher to yield a reliability estimate of the publisher as a whole. This is a relatively well-studied problem. At present, this is usually achieved through a mix of content-based Rashkin et al. (2017); Bianchi et al. (2024) and metadata features Baly et al. (2018a, b, 2019, 2020a, 2020b); Nakov et al. (2021).
Relative to article-level misinformation classifiers, publisher-level classification can greatly reduce the computational cost needed for classification Burdisso et al. (2024). However, this typically involves incorporating additional historical context, world-knowledge Yang and Menczer (2024) or social context Pratelli, Saracco, and Petrocchi (2024). This can make publisher reliability models transductive instead of inductive learners—moderation decisions come from specific prior experience rather than general rules.
This mimics how moderators or users might analyze the reliability of a publisher, potentially before ingesting the contents of a specific article. However, such approaches might fail in cases of where publishers are unknown, ambiguous or evolving. In those situations, moderation decisions at the article-level is necessary. While misinfo-general is suited for either approach, we focus on testing the generalization of inductive article-level classifiers. These models naturally provide classification in cases where limited context or prior experience is available, and are required to depend entirely on specific (and hopefully non-spurious) general rules.
3 Biases in Misinformation Datasets
At risk of repetition: misinformation models’ performance degrades quickly under covariate distribution shifts expected to occur during model deployment, an observation whose cause we attribute to the datasets they were trained on. Due to the exorbitant cost of acquiring high-fidelity misinformation labels, misinformation datasets tend not to reflect the true variance in online (misinformation) content.
To illustrate this, we analyze common properties of datasets specifically constructed for the development of misinformation detectors, by ways of an inexhaustive, yet representative survey of existing misinformation datasets. We provide an overview of these datasets in Table LABEL:tab:misinfo_datasets. Furthermore, in this section, we (1) broadly categorize datasets into different labelling methods; (2) provide specific examples of how misinformation is collected and labelled; (3) discuss how these operationalizations can lead to biases in the datasets; (4) and finally, provide a discussion on the merits and demerits of publisher-level labelled datasets for the purposes of model generalization.
3.1 Dataset Labelling Granularity
Generally speaking, one can classify misinformation datasets into $3$ annotation schemes. Listed from most fine-grained to most coarse-grained:
@@ 3.1 Dataset Labelling Granularity
    Claim: experts fact-check individual (but complete) statements in isolation. Claims are usually small spans sourced from larger documents or utterances
@@
    Article: experts label the overall veracity of entire documents. These can contain many claims, whose factuality need not be consistent with each other
@@
    Publisher: experts label publishers for their propensity for factual reporting, based on historical records and prescribed authorial intent. These labels are often used as a proxy for finer-grained labels. The articles produced by publishers do not necessarily have the same label as the publisher
The more fine-grained annotation methods yield high-quality labels, but can be prohibitively expensive to procure, or evaluate texts without the context those texts would naturally have. Furthermore, these labelling methods are often forced to exclude unverifiable texts (e.g., highly subjective texts or opinions), despite these being prevalent in online discourse. On the other hand, the more coarse-grained annotation methods run the risk of introducing noise into the labels, by assuming consistency between finer-grained labels. For example, an article may contain many factual statements, but a single blatant lie. Since there are increasingly fewer units at each level, however, labels are far easier to procure.
3.2 Survey of Misinformation Datasets
In Table LABEL:tab:misinfo_datasets we present various misinformation datasets, their labelling granularity, their size, and a description of how their data was sampled. In this subsection, we briefly expand on some common trends on how misinformation data and labels were sourced.
‘Claim’-level annotations represent some of the oldest (Lie Detector Mihalcea and Strapparava (2009)) and largest (CREDBANK Mitra and Gilbert (2015)) collections of misinformation text. The claims can be sourced from directly sampling social media (CREDBANK Mitra and Gilbert (2015)) or sampling specific utterances flagged for review (Liar Wang (2017), PolitiFact-Oslo Poldvere, Uddin, and Thomas (2023)).
While labels sourced from domain experts are dominant, using lay people as a method of crowdsourcing for either data or label collection has also proven popular. For the former, as an example, articles are collected only if these were flagged by (trusted) users of social media sites (Weibo15 Ma et al. , Weibo17 Jin et al. (2017), WeChat Wang et al. (2020)). In some cases, lay volunteers were even used in the production of misinformation (Lie Detector Mihalcea and Strapparava (2009), FakeNewsAMT Pérez-Rosas et al. (2018)).
The benefit of crowdsourcing is clear; especially at the article level, datasets that use expert annotations (BuzzFeed-Webis Potthast et al. (2017), Allcott & Gentzkow Allcott and Gentzkow (2017), FakeNewsCorpus Pathak and Srihari (2019)) tend to be much smaller than those leveraging crowdsourcing. A common strategy to combat this, is to blend the ‘Article’ and ‘Publisher’ level labelling schemes (FakeNewsNet/GossipCop Shu et al. (2019), textscFakeNewsCorpus Pathak and Srihari (2019), MM-COVID Li et al. (2020)). Either factual or misinformation articles are manually verified, and the complementary class is sampled from a set of publishers commonly associated with misinformation or factual articles, respectively.
A similar strategy is to blend the ‘Article’ and ‘Claim’ level labelling schemes. A claim made in an article is annotated for veracity in annotation, and its label is propagated to the entirety of the article (FakeNewsNet/PolitiFact Shu et al. (2019), CoAID Cui and Lee (2020), PolitiFact-Oslo Poldvere, Uddin, and Thomas (2023)).
The most consistent method for generating large, diverse corpora, however, proves to be using ‘Publisher-level labelling (TSHP-17 Rashkin et al. (2017), Kaggle Fake News Risdal (2016), Some Like it Hoax Tacchini et al. (2017), Fake vs Satire Golbeck et al. (2018), QProp Barrón-Cedeño et al. (2019)), as discussed above.
Typically, the topics covered in the corpus are not further analyzed by dataset authors, although some datasets specifically focus on articles from various perspectives on the same events (MediaEval15 Boididou et al. , PHEME Zubiaga, Liakata, and Procter (2017), BuzzFeed-Webis Potthast et al. (2017)).
Similarly, while most datasets are fairly general, some focus on specific domains. Very common are those focusing on social media or microblogging texts (CREDBANK Mitra and Gilbert (2015), MediaEval15 Boididou et al. , Weibo15 Ma et al. , Weibo17 Jin et al. (2017), WeChat Wang et al. (2020)). Another common domain involves celebrity rumors, typically annotated for verification rather than veracity, and also commonly sourced from social media posts (Web Dataset Celebrity Pérez-Rosas et al. (2018), FakeNewsNet/GossipCop Shu et al. (2019)). During the COVID-19 pandemic, various health-related datasets were introduced (FakeHealth Dai, Sun, and Wang (2020), MM-COVID Li et al. (2020), FakeCovid Shahi and Nandini (2020), CoAID Cui and Lee (2020)).
3.3 Sources of Dataset Bias
In this subsection, we discuss how specific operationalizations can introduce bias in the dataset, adversely affecting model generalization performance.
Differing Definitions
Even among domain experts, there exists substantial disagreement on what does and does not constitute misinformation Altay et al. (2023). Recent systematic reviews have found that this disagreement has carried over to the computer sciences (see for example Wu et al. (2019); Oshikawa, Qian, and Wang (2020); Zhou and Zafarani (2020); Aïmeur, Amri, and Brassard (2023); Bodaghi et al. (2024); Xiao and Mayer (2024)). Indeed, the surveyed definitions of misinformation in Table LABEL:tab:misinfo_datasets seem to agree on basic properties of misinformation, but disagree on the specific forms. As a result, the forms of misinformation which are included can vary considerably. For example, misinformation forms like ’Satire’ and ’Propaganda’ are either explicitly included or excluded, proving to be especially divisive.
Inconsistent Label Sourcing
Another source of between-dataset variation, is the source of misinformation labels. While most datasets rely on domain experts, some use lay volunteers to verify content, either explicitly (CREDBANK Mitra and Gilbert (2015), Weibo15 Ma et al. , Weibo17 Jin et al. (2017)) or implicitly (Some Like it Hoax Tacchini et al. (2017)).
Recently, datasets have started using many misinformation sources (FakeCOVID Shahi and Nandini (2020), MuMIN Nielsen and McConville (2022), MCFEND Li et al. (2024)). These can come from different countries and cultures, some of which are likely to disagree on their misinformation definitions. Furthermore, this requires aggregating the different misinformation labelling formats.
Most misinformation definitions require specific authorial intent to deceive. However, in some datasets this is missing in the original content (Lie Detector Mihalcea and Strapparava (2009), FakeNewsAMT Pérez-Rosas et al. (2018)), or ambiguous due to misinformation being defined as a lack of credible information (FakeHealth Dai, Sun, and Wang (2020), PHEME Zubiaga, Liakata, and Procter (2017), FakeNewsNet/GossipCop Shu et al. (2019)).
Few publishers
Many datasets limit the number of publishers in either class. In some cases, this is due to deliberate scoping of the dataset (BuzzFeed-Webis Potthast et al. (2017), FakeNewsCorpus Pathak and Srihari (2019)), however in most cases this is due to publisher scarcity. Misinformation annotators, like Snopes, Politifact, GossipCop, etc., understandably tend to focus on verifiable misinformation pieces. As a result, datasets sampling annotations from these sources incur a large positive bias. A common strategy to counteract this is by including samples from a few mainstream publishers (TSHP-17 Rashkin et al. (2017), MM-COVID Li et al. (2020), CoAID Cui and Lee (2020), FakeNewsNet Shu et al. (2019)).
An unwanted side effect of having a small, homogeneous publisher set, is the introduction of a modelling shortcut; misinformation classifiers no longer need to analyze the veracity or intent of input content, but rather simply discriminate between a few publishers with unique idiosyncrasies. Similarly, in datasets where misinformation is constructed by editing factual information (Lie Detector Mihalcea and Strapparava (2009), FakeNewsAMT Pérez-Rosas et al. (2018)), the labels can be inferred by discriminating between the stylistic preferences of the original texts’ authors and those of the editors.
Few events or topics
Similarly, many datasets sample content from a narrow time-span, or from a small set of events or topics. This can reduce the cost of generating labels, but will likely induce overfit in automated moderation systems trained on these corpora.
Focus on Obvious or Popular Misinformation
In several of the discussed datasets, misinformation texts are collected based on user reports, or from third-party fact-checkers. These run the risk of introducing a selection bias, resulting in a dataset that is not representative of all produced misinformation.
A secondary effect of this, is that unverifiable content (e.g., those relying purely on opinion and speculation) are implicitly excluded. Some datasets explicitly exclude unverifiable content (FakeNewsNet Shu et al. (2019)), whereas others include this as a specific category (BuzzFeed-Webis Potthast et al. (2017)). Most datasets do not discuss unverifiable cases, despite these forming a sizable part of produced online content (see Section 8.3).
Conclusion
In short, we find that the realities of misinformation data collection results in many datasets making a trade-off between label quality and corpus size. As a result, these datasets introduce some bias, which we suggest as a primary reason for the reported brittleness of misinformation detectors under covariate shift. Given that these covariate shifts are practically guaranteed in online content or news, testing misinformation detectors before deployment for generalizability is crucial. Doing so, however, requires large, diverse datasets, which we have established is difficult to procure without bias. A related task, publisher reliability estimation, might provide an alternative.
3.4 Publisher Reliability Datasets
Related to the task of misinformation detection is publisher reliability estimation (see Section 2.2. Given an article, or a set of articles, from some publisher, a reliability estimator has to predict the overall publisher reliability.
Publisher reliability is a broader concept than factuality, and considers many aspects of a publisher, which are not necessarily clear when analyzing articles or claims from a publisher in isolation. These aspects include framing, publisher political or editorial bias, intended audience, sourcing practices, funding, etc. All of these factors are analyzed on a large collection of a publishers’ works, and used to provide an indication of the trustworthiness of past and future releases. Ultimately, however, the factuality of produced articles is an important dimension of publisher reliability.
Much like the publisher-level misinformation labelling scheme discussed above, it does not preclude less reliable publishers producing reliable content, or vice versa. It merely suggests that this is less likely to occur. Reliable publishers often produce sensationalist or subjective content to draw in readership, whereas unreliable publishers might intersperse their less reliable articles with more reliable ones to boost their perceived trustworthiness.
Implicitly, by using publisher-level labels as a proxy for article-level reliability, we (as well as many ’Publisher’-level datasets) make the assumption that the article-level factuality of an article from a reliable publisher is stochastically higher than that of an article from a less reliable publisher.
3.4.1 Measuring Generalization with Publisher Reliability Labels
Relative to misinformation datasets, for generalizability aspects, publisher-reliability datasets are far easier to produce at scale, and given publisher-level metadata, can be built specifically to enforce diversity in both publishers and text. Furthermore, articles can be collected across much longer time-spans, which naturally includes shifts in article topics or events.
Perhaps most importantly, however, if a large enough set of publishers is collected, the resulting dataset becomes a naturalistic view of published online content and (mis)information. Instead of a dataset including only verified misinformation, which are typically the least ambiguous or popular cases due to the selection bias of third-party fact-checkers, the dataset is more aligned with online content as it would appear post deployment (Section 8). As a result, statistics about model evaluation are more representative, and model developers can derive stronger conclusions.
In this article, we propose using a publisher-level reliability estimation dataset for the purpose of evaluating the generalizability of article-level misinformation detectors. While this runs the risk of tarring all articles from a publisher with the same brush, we believe the size and diversity of the dataset, along with access to publisher-level metadata, can offset the induced bias, and still allow for conclusive inferences about model behavior under distribution shift.
Specifically, we assume that the effect of covariate distributional shifts on the predictive quality of a model is positively correlated between the two labelling approaches. In other words, we assume that model performance degrades under the same distributional shift in both labelling set-ups. Thus, implicitly we assume that the level of robustness to distributional shifts on a dataset like ‘misinfo-general‘ serves as a good indicator for robustness in article-level misinformation detection.
4 The misinfo-general Dataset
To that end, we introduce misinfo-general, a benchmark for testing the generalization capacity of misinformation detection models, built on top of a series of noisy publisher-level datasets. While best suited for publisher reliability estimation models, we instead use the publisher labels as a proxy for article labels.
Based on the prior discussion, we foresee two sources of bias, (1) labels might not be accurate at the article level, and (2) models will learn to infer the article’s publisher and its label instead of inferring the label from the article. We take the following steps to mitigate these biases as much as possible:
  1. 1.
    relabelling existing articles (Section 4.2)
  2. 2.
    masking or removing publisher identifiable text in articles (Section 4.3)
  3. 3.
    removing any article- and sentence-level duplicates (Section 4.3)